# Color Bind: Exploring Color Perception in Text-to-Image Models

### v2

@@ 4.2.2 Getting the right colors
where $\lambda_{CL}$ is a predefined balancing coefficient and $M^{simp}_{object}$ is a segmentation map extracted for each object from an off-the-shelf segmentation model [34]. In the supplementary material, we show that these masks can also be extracted from the cross-attention maps directly.
Finally, we want to ensure that our modifications do not alter the structure of the original image. To this end, we adopt existing strategies from prior diffusion-based editing methods [16, 7, 21, 29]. Specifically, we replace self-attention maps with those from the simplified prompt and apply background blending during the backward process; additional details are provided in the supplementary.
@@ 4.3 Evaluation
Figure 4: Images generated using the baselines, before and after being edited by our approach, AnySD [41], MagicBrush [42], InstructPix2Pix [5], FPE [21] and MasaCtrl [7]. We illustrate close color pair in the left images. Many additional examples are provided in the supplementary material. Our method consistently outperforms existing editing approaches, we correctly edit the colors and maintain the objects’ shape and background color perfectly aligned with the source image.
To assess the effectiveness of our approach, we compare against several representative editing methods: (i) instruction-following editors (MagicBrush [42], InstructPix2Pix [5], AnySD [41]) trained on paired instruction–image edits, and (ii) attention-manipulation editors (MasaCtrl [7], FPE [21]) that steer self/cross-attention at inference. Implementation details for each method are provided in the supplementary, as well as runtime comparisons, ablations of each module (Stroop loss, color loss, simplified prompt), and results over various metrics that quantify additional aspects, such as semantic alignment and image quality. We also provide interactive visualizations and comparisons over random samples in our benchmark.
@@ Appendix B Evaluation Protocol and Metrics
Visual-Question-Answering Score (VQA-Score): We use a recent multimodal score to evaluate the alignment of the edited image with the input prompt. We calculated VQA-Score [20] of the edited and initial images with respect to the original text prompt (”a color1 colored object1 and a color2 colored object2”). High score indicates high similarity between the input prompt and the image. We followed the setup in the original paper and use the highest-performing model clip-flant5-xxl.
Image-Reward. We calculated ImageReward [39] of the edited image with respect to the original text prompt (”a color1 colored object1 and a color2 colored object2”). ImageReward is a learned reward model trained on 137k expert-labeled human preference comparisons, which assigns higher scores to images that are both visually pleasing and semantically aligned with their text prompts.
@@ C.1.2 Editing Procedure
    Querying Intermediate Latents: During the backward diffusion process we follow [7, 21] using the intermidate latents from the inversion process to improve reconstruction fidelity and editing precision.
@@
    • SAM-based segmentation, we used SAM [34] official GitHub repository, In cases where multiple instances of the same object class are present, we merge their segmentations into a single mask.
    • K-Means clustering on self- and cross-attention maps, following [29], with parameters num_segments=8num\_segments=8 and background_segment_threshold=0.35background\_segment\_threshold=0.35.
@@
    Self-Attention Replacement: We followed [21] and perform self-attention replacements for the first 40 steps, replacing the self-attention maps in layers 4-14.
@@
Finally, we want to ensure that our modifications do not alter the structure of the original image. To preserve the structural integrity of the original image, we draw inspiration from prior work in diffusion-based image editing [16, 7, 21] and replace the self-attention for the full prompt with the self-attention obtained using the simplified prompt. In addition, to maintain the background, we perform blending at time $t_{b}$ similar to [29] in order to preserve the background appearance during the editing process (see Figure 3 in the main paper). Concretely, for each object, we use the object masks $M^{simp}_{object}$ introduced previously in Section 4.2.2 in main paper and take the union to obtain $M^{simp}_{fg}$. Then, we perform blending at time $t_{b}$ during the backward process, updating the latent as follows:
@@ C.2.2 Inference-Time Models
    Bounded-Attention: Based on SD 1.5, using default bounding boxes provided by the authors. For single-object prompts, we used the first default bounding box. Additionally Bounded-Attention requires the input prompt to be split into subject token indices. We use the token division tool provided by [33] to generate it.
@@ C.2.3 Post-Generation Editing Methods
In addition to the methods above, we further compare our editing method with AnySD [41], MagicBrush [42], InstructPix2Pix[5], FPE [21] and MasaCtrl [7]. We used the default setup from the official GitHub repository for them. Additionally, for MasaCtrl, we followed their recommendation to inject the reference latents from the inversion process to the backward process for better reconstruction and editing results. To enable multi-object color editing with AnySD. For AnySD and MagicBrush we followed the official instruction prompt from AnySD applying two consecutive single-object edits, one for each color-object pair. For InstructPix2Pix we follow their official editing prompt, applying two consecutive single-object edits, one for each color-object pair as it resulted in the best performance.
@@ D.2 Evaluation Using Different Metrics
To address the limitations of these traditional metrics, we also incorporate VQAScore [20], which considered the SOTA for measuring similarity between image and text. Our editing method outperforms or is comparable to AnySD, MagicBrush, InstructPix2Pix, FPE and MasaCtrl across all model datasets, except for FLUX-close, where the difference is minimal. However, even VQAScore, despite being a more advanced metric, still struggles to fully capture fine-grained semantic alignment, as illustrated in Figure 2 of the main paper. The limitations of existing evaluation metrics further emphasize the importance of our explicit evaluation pipeline, which directly quantifies perceptible semantic differences in color alignment, providing a more reliable assessment of model performance in compositional color tasks.
@@ D.3 Comparison with ColorPeel
We perform an additional comparison to ColorPeel [6], a text-to-image model that uses prompt learning to learn specific RGB colors embeddings. We followed their official GitHub repository and created four colors:<c1*>,<c2*>,<c3*>,<c4*>. We generated images using prompts inspired by their work. As illustrated in Figure S4 we see that even though ColorPeel specifically targeted the issue of correctly learning RGB color representation, it still fails to generate the colors correctly in the multi-color multi-object setting, as it is challenged by issues such as leakage and swapping of colors. We see that our editing method is able to correctly bind the colors to the objects, resulting in edited images that correspond to the input prompt.
@@ D.6 Performance without using SAM masks
As mentioned in the main paper, we use SAM [34] to extract segmentation masks which are utilized in our editing approach. Next we demonstrate that our approach can also operate without these external masks. For this variant, we use K-means based on the self and cross-attention map during the backward process to extract the objects’ masks, similarly to the approach proposed in [30]. As can be seen on Table S8, the overall results are very similar indicating that our method is not highly-dependent on the segmentation masks obtained by SAM.
@@ D.11.1 Propriety Models Experiment
Figure S7: Propriety model performance on the prompt shown in Figure 1 (main paper). As shown, even the latest VLMs, including OpenAI 4O, Gemini 2.5, and Grok 3, often struggle to capture the correct semantics associated with multiple colors. Our method consistently outperforms existing editing approaches such as AnySD [41], MagicBrush [42], InstructPix2Pix [5], FPE [21], and MasaCtrl [7], which exhibit limitations in addressing such semantic misalignment. MagicBrush and InstructPix2Pix process the input image differently resulting scale change compare to other models with respect to grok3 image.

### v3

@@ 4.2.2 Getting the right colors
where $\lambda_{CL}$ is a predefined balancing coefficient and $M^{simp}_{object}$ is a segmentation map extracted for each object from an off-the-shelf segmentation model [ravi2024sam2]. In the supplementary material, we show that these masks can also be extracted from the cross-attention maps directly.
Finally, we want to ensure that our modifications do not alter the structure of the original image. To this end, we adopt existing strategies from prior diffusion-based editing methods [hertz2022prompttopromptimageeditingcross, Cao_2023_ICCV, liu2024understandingcrossselfattentionstable, Patashnik_2023_ICCV]. Specifically, we replace self-attention maps with those from the simplified prompt and apply background blending during the backward process; additional details are provided in the supplementary.
@@ 4.3 Evaluation
Figure 4: Images generated using the baselines, before and after being edited by our approach, AnySD [yu2025anyeditmasteringunifiedhighqualityz], MagicBrush [zhang2024magicbrushmanuallyannotateddataset], InstructPix2Pix [brooks2023instructpix2pixlearningfollowimage], FPE [liu2024understandingcrossselfattentionstable] and MasaCtrl [Cao_2023_ICCV]. We illustrate close color pair in the left images. Many additional examples are provided in the supplementary material. Our method consistently outperforms existing editing approaches, we correctly edit the colors and maintain the objects’ shape and background color perfectly aligned with the source image.
To assess the effectiveness of our approach, we compare against several representative editing methods: (i) instruction-following editors (MagicBrush [zhang2024magicbrushmanuallyannotateddataset], InstructPix2Pix [brooks2023instructpix2pixlearningfollowimage], AnySD [yu2025anyeditmasteringunifiedhighqualityz]) trained on paired instruction–image edits, and (ii) attention-manipulation editors (MasaCtrl [Cao_2023_ICCV], FPE [liu2024understandingcrossselfattentionstable]) that steer self/cross-attention at inference. Implementation details for each method are provided in the supplementary, as well as runtime comparisons, ablations of each module (Stroop loss, color loss, simplified prompt), and results over various metrics that quantify additional aspects, such as semantic alignment and image quality. We also provide interactive visualizations and comparisons over random samples in our benchmark.
@@ Appendix B Evaluation Protocol and Metrics
Visual-Question-Answering Score (VQA-Score): We use a recent multimodal score to evaluate the alignment of the edited image with the input prompt. We calculated VQA-Score [lin2024evaluatingtexttovisualgenerationimagetotext] of the edited and initial images with respect to the original text prompt (”a color1 colored object1 and a color2 colored object2”). High score indicates high similarity between the input prompt and the image. We followed the setup in the original paper and use the highest-performing model clip-flant5-xxl.
Image-Reward. We calculated ImageReward [xu2023imagerewardlearningevaluatinghuman] of the edited image with respect to the original text prompt (”a color1 colored object1 and a color2 colored object2”). ImageReward is a learned reward model trained on 137k expert-labeled human preference comparisons, which assigns higher scores to images that are both visually pleasing and semantically aligned with their text prompts.
@@ C.1.2 Editing Procedure
    Querying Intermediate Latents: During the backward diffusion process we follow [Cao_2023_ICCV, liu2024understandingcrossselfattentionstable] using the intermidate latents from the inversion process to improve reconstruction fidelity and editing precision.
@@
    • SAM-based segmentation, we used SAM [ravi2024sam2] official GitHub repository, In cases where multiple instances of the same object class are present, we merge their segmentations into a single mask.
    • K-Means clustering on self- and cross-attention maps, following [Patashnik_2023_ICCV], with parameters num_segments=8num\_segments=8 and background_segment_threshold=0.35background\_segment\_threshold=0.35.
@@
    Self-Attention Replacement: We followed [liu2024understandingcrossselfattentionstable] and perform self-attention replacements for the first 40 steps, replacing the self-attention maps in layers 4-14.
@@
Finally, we want to ensure that our modifications do not alter the structure of the original image. To preserve the structural integrity of the original image, we draw inspiration from prior work in diffusion-based image editing [hertz2022prompttopromptimageeditingcross, Cao_2023_ICCV, liu2024understandingcrossselfattentionstable] and replace the self-attention for the full prompt with the self-attention obtained using the simplified prompt. In addition, to maintain the background, we perform blending at time $t_{b}$ similar to [Patashnik_2023_ICCV] in order to preserve the background appearance during the editing process (see Figure 3 in the main paper). Concretely, for each object, we use the object masks $M^{simp}_{object}$ introduced previously in Section 4.2.2 in main paper and take the union to obtain $M^{simp}_{fg}$. Then, we perform blending at time $t_{b}$ during the backward process, updating the latent as follows:
@@ C.2.2 Inference-Time Models
    Bounded-Attention: Based on SD 1.5, using default bounding boxes provided by the authors. For single-object prompts, we used the first default bounding box. Additionally Bounded-Attention requires the input prompt to be split into subject token indices. We use the token division tool provided by [NEURIPS2023_0b08d733] to generate it.
@@ C.2.3 Post-Generation Editing Methods
In addition to the methods above, we further compare our editing method with AnySD [yu2025anyeditmasteringunifiedhighqualityz], MagicBrush [zhang2024magicbrushmanuallyannotateddataset], InstructPix2Pix[brooks2023instructpix2pixlearningfollowimage], FPE [liu2024understandingcrossselfattentionstable] and MasaCtrl [Cao_2023_ICCV]. We used the default setup from the official GitHub repository for them. Additionally, for MasaCtrl, we followed their recommendation to inject the reference latents from the inversion process to the backward process for better reconstruction and editing results. To enable multi-object color editing with AnySD. For AnySD and MagicBrush we followed the official instruction prompt from AnySD applying two consecutive single-object edits, one for each color-object pair. For InstructPix2Pix we follow their official editing prompt, applying two consecutive single-object edits, one for each color-object pair as it resulted in the best performance.
@@ D.2 Evaluation Using Different Metrics
To address the limitations of these traditional metrics, we also incorporate VQAScore [lin2024evaluatingtexttovisualgenerationimagetotext], which considered the SOTA for measuring similarity between image and text. Our editing method outperforms or is comparable to AnySD, MagicBrush, InstructPix2Pix, FPE and MasaCtrl across all model datasets, except for FLUX-close, where the difference is minimal. However, even VQAScore, despite being a more advanced metric, still struggles to fully capture fine-grained semantic alignment, as illustrated in Figure 2 of the main paper. The limitations of existing evaluation metrics further emphasize the importance of our explicit evaluation pipeline, which directly quantifies perceptible semantic differences in color alignment, providing a more reliable assessment of model performance in compositional color tasks.
@@ D.3 Comparison with ColorPeel
We perform an additional comparison to ColorPeel [butt2024colorpeelcolorpromptlearning], a text-to-image model that uses prompt learning to learn specific RGB colors embeddings. We followed their official GitHub repository and created four colors:<c1*>,<c2*>,<c3*>,<c4*>. We generated images using prompts inspired by their work. As illustrated in Figure S4 we see that even though ColorPeel specifically targeted the issue of correctly learning RGB color representation, it still fails to generate the colors correctly in the multi-color multi-object setting, as it is challenged by issues such as leakage and swapping of colors. We see that our editing method is able to correctly bind the colors to the objects, resulting in edited images that correspond to the input prompt.
@@ D.6 Performance without using SAM masks
As mentioned in the main paper, we use SAM [ravi2024sam2] to extract segmentation masks which are utilized in our editing approach. Next we demonstrate that our approach can also operate without these external masks. For this variant, we use K-means based on the self and cross-attention map during the backward process to extract the objects’ masks, similarly to the approach proposed in [patashnik2023localizingobjectlevelshapevariations]. As can be seen on Table S8, the overall results are very similar indicating that our method is not highly-dependent on the segmentation masks obtained by SAM.
@@ D.11.1 Propriety Models Experiment
Figure S7: Propriety model performance on the prompt shown in Figure 1 (main paper). As shown, even the latest VLMs, including OpenAI 4O, Gemini 2.5, and Grok 3, often struggle to capture the correct semantics associated with multiple colors. Our method consistently outperforms existing editing approaches such as AnySD [yu2025anyeditmasteringunifiedhighqualityz], MagicBrush [zhang2024magicbrushmanuallyannotateddataset], InstructPix2Pix [brooks2023instructpix2pixlearningfollowimage], FPE [liu2024understandingcrossselfattentionstable], and MasaCtrl [Cao_2023_ICCV], which exhibit limitations in addressing such semantic misalignment. MagicBrush and InstructPix2Pix process the input image differently resulting scale change compare to other models with respect to grok3 image.